# PriVi: Towards A General-Purpose Video Model For Primate Behavior In The Wild

### v1

@@ Abstract
Non-human primates are our closest living relatives, and analyzing their behavior is central to research in cognition, evolution, and conservation. Computer vision could greatly aid this research, but existing methods often rely on human-centric pretrained models and focus on single datasets, which limits generalization. We address this limitation by shifting from a model-centric to a data-centric approach and introduce PriVi, a large-scale primate-centric video pretraining dataset. PriVi contains 424 hours of curated video, combining 174 hours from behavioral research across 11 settings with 250 hours of diverse web-sourced footage, assembled through a scalable data curation pipeline. We pretrain V-JEPA on PriVi to learn primate-specific representations and evaluate it using a lightweight frozen classifier. Across four benchmark datasets – ChimpACT, BaboonLand, PanAf500, and ChimpBehave – our approach consistently outperforms prior work, including fully finetuned baselines, and scales favorably with fewer labels. These results demonstrate that primate-centric pretraining substantially improves data efficiency and generalization, making it a promising approach for low-label applications. Code, models, and the majority of the dataset will be made available.
@@ 1 Introduction
Figure 1: Our preprocessing pipeline for PriVi. YouTube data is filtered using a learned relevance classifier, while R&O is subsampled based on sub-datset metadata. We apply zero-shot primate detection on both to generate bounding boxes and discard empty frames. PriVi consists of 424 h of unique video complete with bounding boxes and CLIP embeddings for keyframes.
@@
Recent foundation models demonstrate impressive generalization capabilities across many computer vision tasks, including – to some extent – animal behavior [51, 61]. Yet, pretraining data for those models is still very human-centric [4]. While there have been recent efforts to curate large-scale diverse animal video datasets [60, 59, 38], these datasets contain few to no primates. As a result, current video foundation models likely fall behind their full potential when analyzing primate behavior across a wide range of datasets with few labels. What is needed is both a large-scale pretraining datasets and a flexible approach for utilizing the foundation model to solve specific tasks on datasets. In this paper, we work towards filling these two gaps. Our contributions are:
@@
  • Our approach outperforms prior work across four primate behavior datasets.
@@ 2 Related Work
Learning useful representations from unlabeled data has been a cornerstone of machine learning research for decades [53, 25, 24, 55, 29, 43]. The recent success of large language models using masked language modeling [15, 40, 10] popularized self-supervised learning (SSL) as we know it today. There exists a plethora of SSL techniques ranging from contrastive learning/deep metric learning [41, 14] over self-distillation [11] to masked reconstruction in pixel [54, 23] or latent space [4, 47, 1, 2], with many modern approached combining several of those paradigms [39, 62].
The key ingredient for today’s foundation models – large-scale models that can be adapted to a wide range of tasks – is combining SSL with massive Web-scale datasets [49, 37, 31]. These image foundation models, most notably CLIP [41] and DINOv2 [39], caused a paradigm shift from handcrafted and task-specific architectures towards models using foundation models as frozen feature extractors for many computer vision tasks [33, 45]. Recently, there has also been considerable advancement in foundation models for video understanding with several architectures proposed based on video-caption pairs like PerceptionEncoder [5] and VideoCLIP [58], masked video autoencoding [54, 18], masked reconstruction in latent space like V-JEPA [4, 2], and combined approaches like VideoPrism [61].
Web-scale datasets made foundation models possible, but work on data-centric learning [21] shows that neither model nor data scale is sufficient: Careful curation of training data considerably improves performance. Many foundation models thus employ data engines for automated data curation [39, 5, 42, 30] .
@@
There has been considerable prior work for automated behavior recognition: Bain et al. [3] showed that they can discriminate two distinctive actions in wild chimpanzees using audiovisual input. Recently, several methods for more challenging behavior recognition, like distinguishing between all classes in PanAf [8] or ChimpACT [34] have been proposed [35, 7, 6]. However, each of these works focuses on a single dataset and builds on models pretrained on human-centric datasets, like Kinetics [27]. They also rely on considerable amounts of training data, which is costly to acquire.
@@
In recent years, foundation models emerged as a promising feature extractor for behavior recognition. Previous works found that frozen evaluation of VideoPRISM models is state-of-the-art or competitive with specialized methods on ChimpACT [34], KABR [28] – a dataset of Kenyan wildlife –, as well as various datasets of lab rodents [51, 61]. Mamooler et al. [36] show that jointly finetuning a large vision-language model on multiple animal behavior datasets improves performance across all, suggesting a promising step towards unified, general-purpose models without the need for specialized architectures.
@@ 3 Methodology
Our approach combines domain-related self-supervised pretraining with general-purpose foundation models: Instead of training only on a specific primate dataset with limited data to learn representations for this specific task, we pretrain on a broad, large-scale primate dataset to learn representations useful for several tasks. We first describe our pretraining dataset PriVi and the data pipeline used to curate it. Then we describe our framework for primate behavior recognition using broad domain-related pretraining. It consists of the self-supervised training architecture and the frozen evaluation protocol.
3.1 PriVi: Primate-Specific Pretraining Dataset
Research and Observational Data (R&O)
YouTube Data
The R&O datasets features realistic settings close to the setup of existing and future applications of primate behavior analysis. However, it still has low diversity compared to web-scale pretraining datasets. To overcome this limitation, we scrape a large dataset of YouTube videos to increase diversity. We search YouTube for playlists of videos of primates in the wild and manually select 19 playlists with long, high-quality videos. Downloading these playlists yielded a corpus of 458 h of raw video material. We find that YouTube videos are diverse but on average low quality, with many frames containing e. g. interviews with human researchers, primates in cities or otherwise undesired content. We thus apply a data filtering and curation pipeline (see Section 3.2 below).
Dataset versions: PriVi, YT-Filtered, R&O, and YT-Random
After filtering, we obtain 250 unique hours of curated YouTube videos (which we refer to as YT-Filtered) and 174 unique hours of curated research and observational data (which we refer to as R&O). Both together comprise our pretraining dataset PriVi. It consists of 720,000 three-second video snippets, which are partially overlapping, depending on the source dataset. 63 % of the snippets are from YT-Filtered and 37 % from R&O. For each video snippet, the dataset contains bounding boxes and clip embeddings for the center frame. For R&O, we provide species labels inferred from the source-dataset-specific species information and a zero-shot object detector (see Section 3.2).
We release PriVi except for one source dataset comprising 14 % of our pretraining data. This source dataset contains camera trap footage from ongoing research projects and cannot yet be shared publicly. Consult Table 1 for a breakdown of the distribution of species and settings in the dataset. For ablation purposes, we create YT-Random, a YouTube dataset of the same size as YT-Filtered, but with a random selection of videos instead of the curated set.
@@ 3.2 Data Curation Pipeline
Previous work [21, 39] has repeatedly shown that even for web-scale datasets, good data curation is crucial for model performance. On the scale of hundreds of hours of video material, manual curation is infeasible. We thus automatically filter the dataset using a relevance scoring model on latent representations and zero-shot object detectors (see Fig.˜1 for an overview).
@@
Unlike naturalistic videos, YouTube videos feature frequent cuts. Having cuts in our training videos is undesirable, as the scene before and after the cut might be completely different, making self-supervised modeling harder and less aligned with cut-free downstream datasets. Following [60], we detect cuts in the video by using an off-the-shelf cut detector [12]. Even after cut detection, a longer scene might contain both relevant and irrelevant parts, e. g. a primate leaving the frame after half the clip. To select only relevant video snippets, we chunk videos in 3-s snippets with a stride of 2 s, allowing for an 0.5 s overlap with the previous and next snippet each, and determine for each snippet whether it is relevant for training. We specify a list of inclusion criteria to filter only non-computer-generated videos prominently featuring primates. To allow for scalable relevance filtering, one annotator manually labeled 2,500 images based on these criteria. Automated decision of these inclusion criteria requires image-level summary information, so we train a classifier model on top of CLIP embeddings of center frames for relevance filtering. Our 2-layer MLP classifier achieves 82.8 % recall and 90.3 % precision on a held-out validation set with the threshold being optimized for higher precision than recall (ROC-AUC 95.9 %).
@@ 3.3 Self-Supervised Pretraining
Figure 2: Our pretraining and evaluation architecture. EMA: exponential moving average. Refer to caption trainable parameters. Refer to caption frozen parameters. A. Continual pretraining of V-JEPA [4] using PriVi. B. We perform frozen evaluation with a deep, but narrow classifier. Images from our dataset and PanAf500 [8]
We utilize the V-JEPA architecture [4], a self-supervised learning approach that does not require video–text pairs because it relies on masked reconstruction in latent space. We use its default settings.
The V-JEPA architecture consists of a context encoder $E$, a target encoder $\overline{E}$, and a predictor $P$ (see Fig.˜2). The input is a video tensor $x\in\mathbb{R}^{16\times 224\times 224}$ consisting of 16 frames and with a resolution of $224\times 224$ pixels. The input tensor is obtained by performing random cropping around the bounding boxes predicted by the zero-shot primate detector (Section 3.2). The input is tokenized by using non-overlapping $2\times 16\times 16$ patches and combined with sine-cosine positional embeddings to produce $N=1568$ tokens $X\in\mathbb{R}^{D\times N}$, with token embedding dimension $D=1024$.
During training, part of the input sequence is masked and the training objective is to predict the latent representation of masked tokens given the unmasked tokens (see Fig.˜2). Let Mask and $\text{Mask}^{C}$ be functions that split a set of tokens into the set of masked and unmasked tokens, respectively. The loss for an unlabeled data sample $X$ is
@@ 3.4 Supervised Attentive Classifier
Frozen evaluation of vision backbones is not only a means to demonstrate feature quality, it is also competitive for action recognition, particularly on small datasets [51, 61]. We aim to design an attentive classifier for primate behavior recognition on a frozen pretrained model. Existing designs are suprisingly large: V-JEPA [4] uses one self-attention block (12 M parameters) and V-JEPA2 [2] three self-attention plus one cross-attention block (49 M). We hypothesize that this is overparametrized for typical small-scale animal behavior datasets, making models prone to overfitting and instable training dynamics
To reduce parameter count while keeping the classifier sufficiently deep to prevent underfitting, we downproject from $D=1024$ to $D^{\prime}=64$ dimensions before the first self-attention block in the head (Figure 2). The small latent dimension might make it difficult for one CLS token to aggregate information about all classes (e. g. 23 classes in ChimpACT). To avoid information bottlenecks, we concatenate $C$ learned CLS tokens (one per class, initialized randomly) to the patch tokens.
In summary, given $N$ patch tokens $x_{i}$ produced by the backbone $\overline{E}$ and $C$ trainable CLS tokens $q$, predicted class probabilities $\hat{y}\in[0,1]^{C}$ are calculated as
@@ 4.1 Datasets and Evaluation Setup
We evaluate our approach on four video datasets of primate behavior: PanAf [8], ChimpBehave [19], BaboonLand [16], and ChimpACT [34].
@@
ChimpACT [34] contains 2 h of video footage of chimpanzees recorded at the Leipzig Zoo. It features frame-wise bounding boxes and multi-label behavior annotations across 23 classes. Two different evaluation protocols exist: one with [34] and one without [35] ground truth bounding boxes. We follow the evaluation protocol utilizing ground truth bounding boxes. We predict the labels for a primate $i$ at frame $j$ by sampling a 64-frame miniclip around frame $j$ and producing a crop centered at $i$’s bounding box with padding to incorporate scene context.
@@
Following the established evaluation protocol [16, 8, 19], we report top-1 accuracy (Acc) and average per-class accuracy (balanced accuracy, B-Acc) for PanAf500, BaboonLand, ChimpBehave. As ChimpACT is a multi-label dataset, it uses mAP instead, following the AVA protocol [22]. In addition to mAP, which weights all classes equally, we also report weighted mean average precision (mAPw), where each class average precision (AP) is weighted by its support. Thus, we report one metric weighting all classes equally (B-Acc, mAP) and one metric weighting classes by their support (Acc, mAPw) for each dataset. BaboonLand and ChimpBehave do not contain a validation split. We therefore report ablations on ChimpACT and PanAf500 only.
4.2 Architecture and Pretraining Details
Pretraining
Following V-JEPA [4], we choose a ViT-L model with video input for $E$ and a narrow 12-layer transformer for the V-JEPA predictor $P$. $E$ has 304 M parameters, $P$ has 22 M parameters. We initialize the weights from a checkpoint pretrained on VideoMix2M, consisting of Kinetics, SomethingSomethingv2, and HowTo100M, and perform pretraining for 75 k steps with a batch size of 80. This corresponds to eight epochs on the PriVi dataset. Following Singh et al. [50], we perform warmup and train with a constant small learning rate of $1.5\text{\times}{10}^{-5}$. Training takes around 11 hours on a single node with 4 A100 GPUs.
Attentive Classifier
On ChimpACT and PanAf500, we report performance at the best validation accuracy, for ChimpBehave and BabboonLand, we report performance at the end of training. To reduce noise from small datasets, we always calculate class scores on three views per test sample and train five attentive classifiers and report the average. We always sample 16 frames uniformly from the input miniclip. We use cross-entropy loss except for BaboonLand where standard protocol is EQL loss [52].
5 Results
Table 2: All parts of the PriVi dataset improve behavior recognition. We ablate our dataset composition. We compare pretraining on only human-centric data (VideoMix2M, V-JEPA’s default training data) with in-domain pretraining, pretraining on uncurated YouTube videos (YT-Random) and our dataset. All results are obtained on validation sets using our frozen attentive classifier, see Fig.˜2. Cross-domain comparisons in gray.
@@ 5.1 PriVi improves performance over human-centric and in-domain pretraining
We start our analysis by exploring how much PriVi improves action recognition for a fixed, frozen backbone and attentive readout architecture. We initialize using V-JEPA weights pretrained on VideoMix2M and then pretrain with only YT-Filtered, only R&O and both combined (i. e. PriVi). Both YT-Filtered and R&O individually already improve performance on PanAf and ChimpACT over standard V-JEPA, even though R&O shows smaller gains (Table˜2). R&O contains camera trap recordings of Chimpanzees, but no recordings of zoo-housed chimpanzees, which is mirrored in the larger performance improvements on PanAf500 compared to ChimpACT. Combining YT-Filtered and R&O to PriVi yields further gains on all metrics (Table˜2), underscoring the importance of all parts of the dataset.
Table 3: In-Domain continual pretraining further improves PriVi performance We show continual pretraining on ChimpACT and PanAf500 without labels after pretraining on PriVi. All results are obtained on validation sets using our frozen attentive classifier, see Fig.˜2. Cross-domain comparisons in gray.
Continual Pretrain Data Unique Hours ChimpACT PanAf500
mAP mAPw Acc B-Acc
PriVi $424.0$ $38.75$ $54.32$ $89.65$ $79.95$
+ ChimpACT $1.4$ 41.43 57.22 84.93 69.05
+ PanAf500 $1.5$ 32.90 48.91 90.53 87.29
Table 4: Our method outperforms prior methods. We compare V-JEPA pretrained on PriVi without and with continual in-domain pretraining (CID) on the target dataset to various baseline methods including the current state-of-the-art on each dataset. AlphaChimp solves a harder task by simultaneously predicting bounding boxes; results are our own reproduction due to incompatible evaluation protocols, see Appendix.
PanAf500 ChimpACT BaboonLand ChimpBehave
Acc B-Acc mAP mAPw Acc B-Acc Acc B-Acc
X3D [17] $80.00$ $50.35$ $27.05$ $51.60$ $64.89$ $31.41$ $89.3$ $62.8$
V-JEPA [4] $81.96$ $60.19$ $30.85$ $52.89$ $72.14$ 34.57 $94.62$ $69.06$

### v2

@@ Abstract
Non-human primates are our closest living relatives, and analyzing their behavior is central to research in cognition, evolution, and conservation. Computer vision could greatly aid this research, but existing methods often rely on human-centric pretrained models and focus on single datasets, which limits generalization. We address this limitation by shifting from a model-centric to a data-centric approach and introduce PriVi, a large-scale primate-centric video pretraining dataset. PriVi contains 424 hours of curated video, combining 174 hours from behavioral research across 11 settings with 250 hours of diverse web-sourced footage, assembled through a scalable data curation pipeline. We continue pretraining V-JEPA, a large-scale video model, on PriVi to learn primate-specific representations and evaluate it using a lightweight frozen classifier. Across four benchmark datasets – ChimpACT, PanAf500, BaboonLand, and ChimpBehave – our approach consistently outperforms prior work, including fully finetuned baselines, and scales favorably with fewer labels. These results demonstrate that primate-centric pretraining substantially improves data efficiency and generalization, making it a promising approach for low-label applications. Code, models, and the majority of the dataset will be made available.
Figure 1: Continual pretraining on our diverse, large-scale dataset PriVi surpasses state-of-the-art (SOTA) models across four behavior recognition datasets. We recognize behaviors using frozen evaluation. Continual in-domain pretraining (CID) using self-supervised learning (SSL) further improves performance on most datasets. We show relative improvement of mAP (ChimpACT) or accuracy (others) compared to SOTA. Images: ours, [8].
@@ 1 Introduction
Figure 2: Our preprocessing pipeline for PriVi. YouTube data is filtered using a learned relevance classifier, while R&O is subsampled based on source dataset metadata. We apply zero-shot primate detection on both to generate bounding boxes and discard empty frames. PriVi consists of 424 h of unique video complete with bounding boxes and CLIP embeddings for keyframes.
@@
Recent foundation models generalize impressively across many computer vision tasks, including – to some extent – animal behavior [53, 63]. Yet, pretraining data for those models is still very human-centric [4]. While there have been recent efforts to curate large-scale diverse animal video datasets [62, 61, 40], these datasets contain few to no primates. As a result, current video foundation models likely fall behind their full potential when analyzing primate behavior across a wide range of datasets with few labels. What is needed is both a large-scale pretraining datasets and a flexible approach for utilizing the foundation model to solve specific tasks on datasets. In this paper, we work towards filling these two gaps. Our contributions are:
@@
  • Our approach outperforms prior work across four primate behavior datasets. (Fig.˜1).
@@ 2 Related Work
Unsupervised representation learning has been a cornerstone of machine learning research [55, 26, 25, 57, 30, 45]. The recent success of large language models using masked language modeling popularized self-supervised learning (SSL) as we know it today [15, 42, 10]. A plethora of SSL techniques exists, ranging from contrastive and deep metric learning [43, 14] over self-distillation [11] to masked reconstruction in pixel [56, 24] or latent space [4, 49, 1, 2], with many modern approaches combining several of those paradigms [41, 64].
The key ingredient for today’s foundation models is combining SSL with massive Web-scale datasets [51, 39, 32]. These image foundation models, most notably CLIP [43] and DINOv2 [41], caused a paradigm shift from handcrafted and task-specific architectures towards models using foundation models as frozen feature extractors for many computer vision tasks [35, 47]. Recently, there has also been considerable advancement in foundation models for video understanding with several architectures proposed based on video-caption pairs like PerceptionEncoder [5] and VideoCLIP [60], masked video autoencoding [56, 18], masked reconstruction in latent space like V-JEPA [4, 2], and combined approaches like VideoPrism [63].
Recent work on data-centric learning [22] shows that neither model nor data scale is sufficient: Careful curation of training data considerably improves performance. Many foundation models thus employ data engines for automated data curation [41, 5, 44, 31] .
@@
There has been considerable prior work for automated behavior recognition: Bain et al. [3] showed that they can discriminate two distinctive actions in wild chimpanzees using audiovisual input. Several methods for more challenging behavior recognition, like distinguishing between all classes in PanAf [8] or ChimpACT [36], have been proposed [37, 7, 6, 48, 9, 19]. However, each of these works focuses on a single dataset and most build on models pretrained on human-centric datasets, like Kinetics [28].
@@
Foundation models are promising feature extractors for behavior recognition. Frozen evaluation of VideoPRISM models is state-of-the-art or competitive with specialized methods on ChimpACT [36], KABR [29] – a dataset of Kenyan wildlife –, as well as various datasets of lab rodents [53, 63]. Jointly finetuning a large vision-language model on multiple animal behavior datasets improves performance across all [38], suggesting a promising step towards unified, general-purpose models without the need for specialized architectures.
@@ 3 Methodology
Our approach combines domain-related self-supervised pretraining with general-purpose foundation models: Instead of training only on a small primate dataset to learn representations for this specific task, we pretrain on a broad, large-scale primate dataset to learn representations useful for several tasks. We first describe our pretraining dataset PriVi and our data curation pipeline. Then we describe our framework for primate behavior recognition using broad domain-related pretraining, which consists of self-supervised training and a frozen evaluation protocol.
3.1 PriVi: Primate-Specific Pretraining Dataset
Research and Observational Data (R&O)
YouTube Data
The R&O dataset features realistic settings close to the setup of existing and future applications of primate behavior analysis. However, it still has low diversity compared to web-scale pretraining datasets. To overcome this limitation, we scrape a large dataset of YouTube videos to increase diversity. We search YouTube for playlists of videos of primates in the wild and manually select 19 playlists with long, high-quality videos. Downloading these playlists yielded a corpus of 458 h of raw video material. We find that YouTube videos are diverse but on average low quality, with many frames containing e. g. interviews with human researchers, primates in cities or otherwise undesired content. We thus apply a data filtering and curation pipeline (see Sec.˜3.2 below).
Dataset versions: PriVi, YT-Filtered, R&O, and YT-Random
After filtering, we obtain 250 unique hours of curated YouTube videos (which we refer to as YT-Filtered) and 174 unique hours of curated research and observational data (which we refer to as R&O). Both together comprise our pretraining dataset PriVi. It consists of 720,000 three-second video snippets, which are partially overlapping, depending on the source dataset. 63 % of the snippets are from YT-Filtered and 37 % from R&O. For each video snippet, the dataset contains bounding boxes and clip embeddings for the center frame. For R&O, we provide species labels inferred from the source-dataset-specific species information and a zero-shot object detector (see Sec.˜3.2).
We release PriVi except for one source dataset comprising 14 % of our pretraining data. This source dataset contains camera trap footage from ongoing research projects and cannot yet be shared publicly. Consult Tab.˜1 for a breakdown of the distribution of species and settings in the dataset. For ablation purposes, we create YT-Random, a YouTube dataset of the same size as YT-Filtered, but with a random selection of videos instead of the curated set.
@@ 3.2 Data Curation Pipeline
Previous work [22, 41] has repeatedly shown that even for web-scale datasets, good data curation is crucial for model performance. On the scale of hundreds of hours of video material, manual curation is infeasible. We thus automatically filter the dataset using a relevance scoring model on latent representations and zero-shot object detectors (see Fig.˜2 for an overview).
@@
Unlike naturalistic videos, YouTube videos feature frequent cuts. Having cuts in our training videos is undesirable, as the scene before and after the cut might be completely different, making self-supervised modeling harder and less aligned with cut-free target datasets. Following [62], we detect cuts in the video by using an off-the-shelf cut detector [12]. Even after cut detection, a longer scene might contain both relevant and irrelevant parts, e. g. a primate leaving the frame after half the clip. To select only relevant video snippets, we chunk videos in 3-s snippets with a stride of 2 s, allowing for an 0.5 s overlap with the previous and next snippet each, and determine for each snippet whether it is relevant for training. We specify a list of inclusion criteria to filter only real-world videos prominently featuring primates. To allow for scalable relevance filtering, one annotator manually labeled 2,500 images based on these criteria. Automated decision of these inclusion criteria requires image-level summary information, so we train a classifier on top of CLIP embeddings of center frames for relevance filtering. Our 2-layer MLP classifier achieves 82.8 % recall and 90.3 % precision on a held-out validation set with the threshold being optimized for higher precision than recall (ROC-AUC 95.9 %).
@@ 3.3 Self-Supervised Pretraining
Figure 3: Our pretraining and evaluation architecture. EMA: exponential moving average. $\|$: sequence concatenation. A. We continue self-supervised pretraining of a V-JEPA model on our primate dataset PriVi, doing masked prediction in latent space. B. We train a classifier for each target dataset. Images are ours or [8].
Figure 4: Example predictions of our model on PanAf500 and ChimpACT. We show annotated ground truth (gt) and predictions (pred) of the model pretrained on PriVi only. Examples from the validation sets. Image sources: [36, 8].
We utilize V-JEPA [4], a self-supervised learning approach that does not require video–text pairs but does masked prediction in latent space. We use its default settings.
The V-JEPA architecture consists of a context encoder $E$, a target encoder $\overline{E}$, and a predictor $P$ (see Fig.˜3). The input is a video tensor $x\in\mathbb{R}^{16\times 224\times 224}$ consisting of 16 frames and with a resolution of $224\times 224$ pixels. The input tensor is obtained by performing random cropping around the bounding boxes predicted by the zero-shot primate detector (Sec.˜3.2). The input is tokenized by using non-overlapping $2\times 16\times 16$ patches and combined with sine-cosine positional embeddings to produce $N=1568$ tokens $X\in\mathbb{R}^{N\times D}$, with token embedding dimension $D=1024$.
During training, part of the input sequence is masked and the training objective is to predict the latent representation of masked tokens given the unmasked tokens (see Fig.˜3). Let Mask and $\text{Mask}^{C}$ be functions that split a set of tokens into the set of masked and unmasked tokens, respectively. The loss for an unlabeled data sample $X$ is
@@ 3.4 Supervised Attentive Classifier
Frozen evaluation of large vision models is not only a means to demonstrate feature quality, it is also competitive for action recognition, particularly on small datasets [53, 63]. We aim to design an attentive classifier for primate behavior recognition on a frozen pretrained model. Existing designs are surprisingly large: V-JEPA [4] uses one self-attention block (12 M parameters) and V-JEPA2 [2] three self-attention plus one cross-attention block (49 M). We hypothesize that this is overparameterized for typical small-scale animal behavior datasets, making models prone to overfitting and unstable training dynamics
To reduce parameter count while keeping the classifier sufficiently deep to prevent underfitting, we downproject from $D=1024$ to $D^{\prime}=64$ dimensions before the first self-attention block in the classifier (Fig.˜3). The small latent dimension might make it difficult for one CLS token to aggregate information about all classes (e. g. 23 classes in ChimpACT). To avoid information bottlenecks, we concatenate $C$ learned CLS tokens (one per class, initialized randomly) to the patch tokens.
In summary, given $N$ patch tokens $x_{i}$ produced by the encoder $\overline{E}$ and $C$ trainable CLS tokens $q_{j}$, predicted class probabilities $\hat{y}\in[0,1]^{C}$ are calculated as
@@ 4.1 Datasets and Evaluation Setup
We evaluate on four primate video datasets: PanAf [8], ChimpBehave [20], BaboonLand [16], ChimpACT [36].
@@
ChimpACT [36] contains 2 h of video footage of chimpanzees recorded at the Leipzig Zoo. It features frame-wise bounding boxes and multi-label behavior annotations across 23 classes. Two different behavior recognition tasks exist: one with [36] and one without [37] access to ground truth bounding boxes. We utilize ground truth boxes. We predict the labels for a primate $i$ at frame $j$ by sampling a 64-frame miniclip around frame $j$ and producing a crop centered at $i$’s bounding box with padding to incorporate scene context.
@@
Following the established evaluation protocol [16, 8, 20], we report top-1 accuracy (Acc) and average per-class accuracy (balanced accuracy, B-Acc) for PanAf500, BaboonLand, ChimpBehave. As ChimpACT is multi-label, it uses mAP instead, following the AVA protocol [23]. In addition to mAP, which weights all classes equally, we also report weighted mean average precision (mAPw), where each class average precision (AP) is weighted by its support. We report mAP and mAPw over 23 classes. Thus, we report one metric weighting all classes equally (B-Acc, mAP) and one metric weighting classes by their support (Acc, mAPw) for each dataset. BaboonLand and ChimpBehave do not contain a validation set. We therefore report ablations on ChimpACT and PanAf500 only.
Table 2: All parts of the PriVi dataset improve behavior recognition. We ablate the pretraining data composition. We compare (a) human-centric data (V-JEPA) and in-domain data only with (b) uncurated YouTube videos (YT-Random), our datasets (PriVi, YT-Filtered, R&O) and continual in-domain pretraining (CID). All experiments on validation sets using our attentive classifier, see Fig.˜3. Cross-domain comparisons in gray.
@@
+ CID: ChimpACT $1.4$ 41.43 57.22 84.93 69.05
+ CID: PanAf500 $1.5$ 32.90 48.91 90.53 87.29
4.2 Architecture and Pretraining Details
Table 3: Our method outperforms prior methods. We compare our attentive classifier with V-JEPA pretrained on only human-centric data (V-JEPA), pretrained on PriVi, and with continual in-domain pretraining (CID) to various baseline and state-of-the-art methods. AlphaChimp solves the harder task of simultaneously predicting bounding boxes and cannot be evaluated with ground truth bounding boxes; results are our own reproduction due to incompatible evaluation protocols, see Appendix. Results on test sets.
ChimpACT PanAf500 BaboonLand ChimpBehave
mAP mAPw Acc B-Acc Acc B-Acc Acc B-Acc
X3D [17] $27.05$ $51.60$ $80.00$ $50.35$ $64.89$ $31.41$ $89.3$ $62.8$